# Learning to accelerate distributed ADMM using graph neural networks

### v1

@@ 6 Discussion & conclusion
In this paper we have seen how the distributed ADMM algorithm can be accelerated by expressing the updates in form of message-passing steps. In the experimental section we utilize this to learn the step size and communication matrix but there are several other ways to further exploit this connection. In the following we highlight some of these potential extensions and discuss limitations of our approach.
@@
With the training scheme outlined in Section 4.2, we train the unrolled algorithm for a fixed number of iterations in our experiments to achieve a low loss. However, ADMM is also known to suffer from slow convergence when getting closer to the minima. With the current training scheme we only train the network to perform well after the $K=10$ unrolled steps which does not necessarily translate into a faster global convergence [4]. Further, we implement the differentiation of the optimization problem using unrolling of the inner solvers which has a larger memory footprint than implicit differentiation and as such limits the number of unrolled ADMM iterations we can train the network on. The specific method is a trade-off between computational time and memory requirements that can be chosen based on the underlying problem class and its dimensionality [35]. However, in practical real-time applications a certain computational budget typically exists, justifying our approach of tuning performance to a fixed number of steps.
@@
In our approach, we only augment distributed ADMM by learning components that predict the hyperparameters. However, it is also conceivable to replace some of the steps completely. For example, the update of the $x$-variable requires solving a – possibly time-consuming – optimization problem. While it is tempting to replace this with the output of a neural network, it is significantly more challenging to ensure convergence for such inexact methods. Nevertheless, it has been successfully demonstrated for other optimization schemes [5]. Further, the exchange of messages involving all iterates can be computationally expensive. Accordingly, the framework can be extended to only communicate compressed information about the iterates without losing convergence [11].
Further, in our paper we focus on the decentralized distributed ADMM version presented by Makhdoumi and Ozdaglar [29] while many other formulations of distributed ADMM exist that can leverage GNNs. In particular GNNs with a global node can be directly mapped to distributed ADMM versions with a centralized communication hub that is connected to all nodes in the graph opening up for many more synergies between distributed optimization and graph neural networks in the future.
Figure 2: Semi-log plots of the relative objective for the two problem classes across all 2020 iterations. The network is only unrolled for the first $10$ iterations during training as in the previous results.
@@ Conclusion
In this paper, we demonstrate an equivalence between distributed ADMM and message-passing networks that, to our knowledge, has not been explicitly recognized before. Based on this insight, we describe a learning-to-optimize approach that predicts instance-specific hyperparameters for distributed ADMM, which is trained by unrolling the algorithm and minimizing the normalized distance between the true minimizer and the network output for a specific problem distribution. These learning tasks can be naturally expressed as different graph learning tasks leveraging the connection between ADMM and GNNs. In numerical experiments, we validate that the learned algorithm outperforms the baseline method and leads to an accelerated convergence of ADMM.
Acknowledgments
The authors thank Daniel Arnström for helpful discussions and suggestions. This research was funded in part by Sweden’s Innovation Agency (Vinnova), grant number 2022-03023, the Centre for Interdisciplinary Mathematics at Uppsala University (CIM), and supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
@@ Appendix B Problem details
For each problem class we generate a total of $900$ instances for training, $100$ instances for validation and $100$ instances for testing purposes. All results in the main text are evaluated on the testing instances which are not seen by the network during training.
@@ B.1 Network average consensus
For the network average consensus problem, each node or agent $i\in\{1,\ldots,m\}$ in the network has a local objective function as in (9) to minimize the distance between the local optimization variable $x_{i}$ and its local private data $b_{i}$.
@@
This local objective function is parameterized for every agent by a vector $b_{i}\in\mathbb{R}^{n}$. It is sampled independently from a $n$-dimensional normal distribution $\mathcal{N}(0,\Sigma)$ with covariance matrix $\Sigma=100I_{n}$. In our experiments, we fix $m=8$ and $n=2$.
@@ B.2 Distributed least squares problem
In the distributed least squares problem, the local objective function is of the form (10). In addition to the data vector $b_{i}$ from the network average consensus problem, each agent also has a private feature transformation matrix $B_{i}\in\mathbb{R}^{n\times n}$. The latter is used to weight the linear regressionc efficients of $x_{i}$ before comparing with $b_{i}$.
@@
Across all problem domains, the underlying graph structure is based on the Erdős-Rényi random graph model [17] with $m=8$ nodes. We discard sampled graphs which are not connected.
@@
Appendix C Training details
In the following, we discuss the details of the training scheme and the chosen hyperparameters in order to ensure reproducibility of our presented results.
Implementation details
We use the JAX framework [9] to implement the differentiable distributed ADMM algorithm as a message-passing network. In particular, we are using the jraph toolkit to implement the message-passing steps [23] and integrate it with the flax library to design neural network components using the novel linen API [27].
@@ Appendix C Training details
  • The input for the edge weight prediction is given by the concatenated local degree profiles of the adjacent nodes [10]. Each of these consists of $5$ features containing statistics about the the neighborhood structure including the degree, maximum and minimum neighboring degree, as well as the mean and variance. To ensure that the communication matrix is symmetric, we sum up the edge predictionsWe use the same predicted communication matrix for all iterations, thus, the number of parameters in the model is $385$.
@@
Loss function
To avoid division by zero, we include the hyperparamter $\varepsilon$ into the loss function in (8). In the experiments, we set $\varepsilon=10^{-5}$.
@@
For the training, we are using the Adam optimizer with a learning rate of $10^{-4}$ and global gradient clipping of with radius $1.0$. We are training the model for $100$ epochs using a batch size of $5$. Thus, during training a total of $18\thinspace 000$ parameter update steps are executed.
@@
All reported experiments are executed on a single NVIDIA TITAN Xp GPU with 12 GB of memory. Training a single epoch requires approximately $10$ minutes plus additional time for the validation process. In total, the training of one model for $100$ epochs requires around $20$ hours.
Appendix D Algorithms
D.1 Decentralized, distributed ADMM
@@
In the following, we state the decentralized, distributed ADMM algorithm (Algorithm 1) from Makhdoumi and Ozdaglar [29] adapted to our notation. As described in section 3.2, each iteration updates all three iterates $(x_{i},y_{i},\lambda_{i})$ and includes two communication steps to pass information about the current iterates to the neighbors in $\mathcal{G}$.
@@
D.2 Message-passing formulation of distributed ADMM

### v2

@@ 6 Discussion & conclusion
We have shown how the distributed ADMM algorithm can be accelerated by expressing the updates in form of message-passing steps that can be parametrized to learn the step size and communication matrix. Several other approaches could further exploit this connection. In the following we highlight some of these potential extensions and discuss limitations of our approach.
@@
ADMM has robust guarantees but is known to suffer from slow convergence when approaching the optimum. Since we intentionally stay within the ADMM framework, we cannot expect to overcome this inherent weakness. We only train the network to perform well within a fixed number of iterations $K$ across all instances. This does not necessarily translate into improved performance on every individual instance, or faster global convergence beyond the $K$ iterations, which is a general limitation of unrolling (arisaka2023principled). We observe this in our experiments, where the combined method leads to a larger error than the baselines when increasing the iterations significantly. However, practical applications typically have a computational budget, making unrolling a useful technique for real-world problems.
@@
We only augment ADMM by learning to predict the hyperparameters. However, it is also possible to replace some of the ADMM steps completely. For example, the update of the $x$-variable requires solving a potentially time-consuming optimization problem (li2023learning). While it is tempting to replace this with the output of a neural network directly, ensuring convergence for such inexact methods is significantly more challenging (banert2024accelerated). Additionally, the exchange of messages involving all iterates can be limited by the network bandwidth. ADMM can be extended to only communicate compressed information without losing convergence (chen2025greedy).
We focus on decentralized distributed ADMM (makh2017admm) while many other formulations of distributed ADMM exist that can leverage graph neural networks. In particular, distributed ADMM with a centralized communication hub can be mapped to message-passing with a global node, opening up many more synergies between distributed optimization and GNNs.
Figure 2: Semi-log plots of the mean relative objective over all $100$ test instances for 20 iterations. The learned methods are only trained on the first $10$ iterations as in previous results.
@@ Conclusion
In this paper, we demonstrate an equivalence between distributed ADMM and message-passing networks that, to our knowledge, has not been explicitly recognized before. Based on this insight, we describe a learning-to-optimize approach that predicts instance-specific hyperparameters for distributed ADMM. The model is trained by unrolling the algorithm and minimizing the normalized distance between the true minimizer and the network output for a specific problem distribution. These learning tasks can be naturally expressed as different graph learning tasks leveraging the connection between ADMM and GNNs. In numerical experiments, we validate that the learned algorithm outperforms the baseline methods and leads to faster convergence of ADMM.
[…]
The authors thank Daniel Arnström for helpful discussions and suggestions and the anonymous reviewers for their thoughtful comments. This research was funded in part by Sweden’s Innovation Agency (Vinnova), grant number 2022-03023, the Centre for Interdisciplinary Mathematics at Uppsala University (CIM), and supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
@@ Appendix A Glossary of Notation
$r,s$ Primal and dual ADMM residuals
$\mu,\tau$ Adaptive step size parameters
@@ Appendix B Problem details
Across all problem domains, the underlying graph structure is based on the Erdős-Rényi random graph model (Erdos:1959:pmd) with $m=8$ nodes. The edge probability is fixed to $p=0.4$ in the experiments. We discard sampled graphs that are not connected. Moreover, the dimensionality of the solution $x^{\star}\in\mathbb{R}^{n}$ is fixed at $n=2$ in all experiments.
For each problem class, we generate a total of $900$ instances for training, $100$ instances for validation and $100$ instances for testing purposes. All results in the main text are evaluated on the testing instances which are not seen by the network during training.
@@ B.1 Network average consensus
For the network average consensus problem, each node or agent $i\in\{1,\ldots,m\}$ in the network has a local objective function that minimizes the distance between the local optimization variable $x_{i}$ and its local private data $b_{i}$.
@@
This local objective function is parameterized for every agent by a vector $b_{i}\in\mathbb{R}^{n}$. The vector is sampled independently from a $n$-dimensional normal distribution $\mathcal{N}(0,\Sigma)$ with covariance matrix $\Sigma=100I_{n}$.
@@ B.2 Distributed least squares problem
In the distributed least squares problem, the local objective function solves the least-squares problem for the privately available data. In addition to the data vector $b_{i}$ from the network average consensus problem, each agent also has a private feature transformation matrix $B_{i}\in\mathbb{R}^{n\times n}$. The latter is used to weight the linear regression coefficients of $x_{i}$ before comparing with $b_{i}$.
@@
Appendix C Implementation details
In the following, we discuss the details of the training scheme and the chosen hyperparameters and tuning steps in order to ensure reproducibility of our presented results.
Framework
We use the JAX framework (jax2018github) to implement the differentiable distributed ADMM algorithm as a message-passing network. In particular, we are using the jraph toolkit to implement the message-passing steps (jraph2020github) and integrate it with the flax library to design neural network components using the novel linen API (flax2020github).
Optimization subproblem
Computing the first part can be achieved in closed form by the means of implicit differentiation (amos2017optnet). However, implicit differentiation can in practice be costly and requires additional overhead. Instead, we solve the subproblem (6) approximately by unrolling an iterative solver [^4] Here, we unroll the iterative solver within each unrolled iteration of ADMM. or use the closed-form solution to (6) if it exists. In both cases, we compute its gradient using standard automatic differentiation techniques. Along the same lines, the partial derivative of the optimization problem (3) with respect to the learned edge weights $e_{ij}$ is computed. The remaining update steps can be differentiated using standard automatic differentiation tools provided by JAX (jax2018github).
@@ Appendix C Implementation details
  • The input for the edge weight prediction is given by the concatenated local degree profiles of the adjacent nodes (cai2018simple). Each of these consists of $5$ features containing statistics about the the neighborhood structure including the degree, maximum and minimum neighboring degree, as well as the mean and variance. To ensure that the communication matrix is symmetric, we sum up the edge predictions. Further, we use the same predicted communication matrix for all iterations. Thus, the number of parameters in the model is $385$.
@@
For the training, we are using the Adam optimizer with a learning rate of $10^{-4}$ and global gradient clipping of with radius $1.0$. We are training the model for $100$ epochs using a batch size of $5$ via gradient accumulation. Thus, during training a total of $18\thinspace 000$ parameter update steps are executed. To avoid division by zero, we include the hyperparameter $\varepsilon$ into the loss function in (7). In the experiments, we set $\varepsilon=10^{-5}$.
ADMM hyperparameters
If not learned or further specified, we apply the naive step size $\alpha=1$ and the Laplacian matrix $\mathcal{L}$ of the graph as the communication matrix $P$ in all algorithm steps. Like the Laplacian, the degree $d_{i}$ of each node $i$ depends on the respective graph instance.
ADMM initialization
In our experiments, we initialize all $x_{i}$ by $\mathbf{0}\in\mathbb{R}^{n}$ and hence also all $y_{i}$ by $\mathbf{0}\in\mathbb{R}^{n}$. While it is generally possible to use warm starting with the proposed framework, e.g. through learning (sambharya2024learning), additional algorithm steps are required in this case to ensure that the consensus variables $y_{i}$ are appropriately initialized. For details, see lines 2-6 in Algorithm 1.
Fixed baseline
The fixed baseline method uses a constant, global $\alpha$ across all iterations, which is chosen via grid search on the validation data. We allow $\alpha$ to be in the interval $[0.001,10]$ using linear spacing for $100$ different values. [^5] We choose the initial interval such that it covers the empirically observed learned step sizes across the unrolled iterations. This is a form of black-box optimization. To choose the $\alpha$ that performs best on average for all instances, we use the same objective as in training, defined in equation (7). For both problem classes, we select the $\alpha$ in our grid that performs best in this metric on the validation data after $K=10$ iterations, and apply it on the test data in the reported results. This leads to selecting $\alpha=1.112$ for the consensus problem, and $\alpha=1.011$ for the least squares problem. This approach is similar to the method proposed by noah2024distributed in the case where only one constant step size is learned, but it relies on black-box optimization rather than gradient-based optimization to determine the parameters.
Adaptive baseline
In the adaptive baseline, the local node-level step sizes $\alpha_{i}$ gets adjusted dynamically based on the balance of primal and dual residuals (boyd2011). Intuitively, if the primal residual is too large, we increase $\alpha$ to shift the focus towards enforcing the consensus, while if the dual residual is too large, we decrease the step size to allow for more flexibility in the local node updates. We utilize the common heuristic introduced by he2000alternating to update the step size in each step as:
$\alpha^{k+1}=\begin{cases}\tau\alpha^{k}\quad&\text{if }\|r^{k}\|>\mu\|s^{k}\|,\\ \tau^{-1}\alpha^{k}\quad&\text{if }\|s^{k}\|>\mu\|r^{k}\|,\\ \alpha^{k}\quad&\text{otherwise}.\end{cases}$ (12)
Here, $\tau\in\mathbb{R}_{+}$ and $\mu\in\mathbb{R}_{+}$ are parameters of the adaptive scheme. Typical default choices are $\tau=2$ and $\mu=10$ (he2000alternating). The primal and dual residual at iteration $k$ are denoted by $r^{k}$ and $s^{k}$ respectively and defined as:
$r_{ij}^{k+1}$ $=P_{ij}x_{j}^{k+1}-z_{ij}^{k+1}=y_{i}^{k+1},$ (13)
$s_{j}^{k+1}$ $=\alpha\sum_{i=1}^{m}\left[P_{ij}^{2}\left(x_{j}^{k+1}-x_{j}^{k}\right)+P_{ij}\left(y_{i}^{k+1}-y_{i}^{k}\right)\right]$
$=\alpha\left(\left[\sum_{i=1}^{m}P_{ij}^{2}\right](x_{j}^{k+1}-x_{j}^{k})+d_{j}(y^{k+1}_{j}-y^{k}_{j})+\bar{y}_{\rightarrow j}^{\,k+1}-\bar{y}_{\rightarrow j}^{\,k}\right)$
The second formulation of the dual residual is readily available in our message-passing framework. Note that the primal residual is an edge-based version of the newly introduced $y^{k}$ in the distributed ADMM framework as defined in equation (4), measuring the constraint violation of the primal problem. The dual residual is a measure of change in this auxiliary variable and the local solution $x_{i}$. In order to compute a node-based primal residual for predicting a node-based $\alpha_{i}$ with the step size heuristic, we use
$\|r_{i}\|_{2}=\sqrt{\sum_{j\in\mathcal{N}(i)}r_{ij}^{T}r_{ij}}=\sqrt{d_{i}\,y_{i}^{T}y_{i}}=\sqrt{d_{i}}\,\|y_{i}\|_{2}.$ (14)
As for the fixed baseline, we choose the hyperparameters of the adaptive scheme using grid search over the validation data, based on average performance after $K=10$ iterations. For the $\mu$ parameter, that controls how frequently the step size parameter is changed, we search over the choices $\{1,5,10,15,20,25,30,25,30,35,40\}$. For the scaling parameter, that adjusts the step size we pick values in the interval $(1,2]$. The grid search picks $20$ different values from the interval with logarithmic spacing. Thus, in total we evaluate $200$ different combinations of parameters for the tuning of the adaptive step size scheme. From this, we obtain the optimal values $\mu=10$ and $\tau=1.32$ for the consensus problem, and similarly $\mu=10$ and $\tau=1.36$ for the least-squares problem. In order to avoid numerical breakdowns, we reset the step size to the fixed value of $\alpha=1$ after the first $K=10$ iterations, as in the other methods.
@@
All reported experiments are executed on a single NVIDIA TITAN Xp GPU with 12 GB of memory. Training a single epoch requires $10$ minutes plus additional time for the validation process. In total, the training of one model for $100$ epochs requires around $20$ hours.
Appendix D Additional results
Here, we provide additional results to complement the empirical evaluation of our methods based on the convergence time, the distribution of solving times, and the out-of-domain (OOD) performances.
Table 3: Timing and convergence comparison of the different methods until the relative objective reaches a threshold $\varepsilon$. The total computational budget for each problem is 11 second. For each solver we check how many instances achieve the desired accuracy within this budget and report the resulting success rate ($\uparrow$ higher is better). For the successfully solved instances per method, we measure the average number of iterations and required time in seconds ($\downarrow$ lower is better). Best is bold, second best underlined.
$\varepsilon=0.05$ $\varepsilon=0.01$ $\varepsilon=0.001$
Consensus Iterations Time Success Iterations Time Success Iterations Time Success
Fixed 11.64 0.23 0.94 16.36 0.30 0.87 29.42 0.50 0.76
Adaptive 11.30 0.30 0.93 15.76 0.36 0.83 26.85 0.57 0.71
Global $\alpha$ 7.82 0.20 0.96 19.37 0.39 0.91 35.31 0.64 0.78
Local $\alpha_{i}$ 5.61 0.16 0.96 12.72 0.28 0.89 26.88 0.51 0.80
Weight $e_{ij}$ 11.19 0.24 0.96 15.62 0.32 0.92 27.38 0.46 0.82
Combined 5.61 0.17 0.98 10.47 0.25 0.96 23.56 0.44 0.90
Least-squares Iterations Time Success Iterations Time Success Iterations Time Success
Fixed 9.61 0.30 0.85 14.72 0.44 0.60 21.23 0.62 0.26
Adaptive 9.20 0.32 0.85 13.29 0.44 0.59 20.23 0.62 0.26
Global $\alpha$ 5.31 0.22 0.94 11.30 0.36 0.76 18.47 0.58 0.34
Local $\alpha_{i}$ 6.27 0.22 0.89 11.67 0.39 0.76 18.97 0.60 0.35
Weight $e_{ij}$ 10.88 0.33 0.86 15.68 0.47 0.63 19.54 0.56 0.39
Combined 5.41 0.21 0.94 10.13 0.34 0.83 17.90 0.53 0.41
D.1 ADMM convergence and solver time
In Table 3, we compare the different methods based on achieving a certain threshold in relative objective value within a given computational budget of one second. For every instance on which a certain method reaches the threshold within the budget, we measure the required number of iterations and wall-clock time. In the table, we report these values as averages across successful instances per method. Further, we report the fraction of successfully solved instances within the budget. Note that this leads to a dependent metric since the iterations and time are conditional on solving. Thus, we need to evaluate the different methods based on their Pareto front instead of looking at the metrics individually. A method is on the Pareto front if no other method achieves both a higher success rate and lower iteration/time performance at the same time.
Across both problem classes, all learned solvers are able to solve more instances within the computational budget than the fixed and adaptive baseline. Moreover, for every threshold, there exist multiple learned methods that are Pareto improvements to the two baselines, i.e. more instances are solved while requiring fewer average iterations and solving them in faster average time. In particular, the combined method appears to be the strongest choice as it is always a Pareto improvement over the baselines, solves the most instances, and is on the Pareto front for all thresholds and problems. Only for $\varepsilon=0.05$ there exists a learned method that requires marginally fewer iterations or less time.
Figure 3: Pairwise comparison of the the distributions of normalized objective value across the test set between the fixed baseline method (gray) and the learned methods (colored). The x-axis shows the log-scaled normalized residual after $10$ iterations.
D.2 Distribution of solving times
In the previous results, we only report the mean performance across all problem instances to summarize the performance in a single number. Here, we additionally show the distribution of the performance across the $100$ individual test instances. For each learned ADMM solver, we compare the distribution of relative objectives after $K=10$ iterations in Figure 3. Overall, we can see that there is a large spread of relative objectives achieved after the iterations showing that the problems are vastly different in difficulty.
One observation is that the learned ADMM method can lead to slightly worse performance for individual instances for which the fixed baseline method works very well. This is due to the fact that the training aims to improve the mean performance across all instances (arisaka2023principled).
D.3 Generalization behavior of learned methods
OOD generalization
To examine the robustness and generalization ability of our learned methods, we evaluate their performance on larger graphs unseen during training, and compare the results with the behavior of the two baselines. For the out-of-distribution samples, we slightly change the data distribution by adjusting the edge weight parameter $p$ in the random graph generation to keep the approximate neighborhood structure fixed. Given the initial value for $p=0.4$ the expected number of neighbors in a graph with $m=8$ nodes is $2.8$. We adjust the edge probability for a graph with $m\in\{16,32,64,128\}$ nodes to p=2.8/(m1)p=2.8/(m-1). Thus, the expected degree of each node remains the same even when increasing the graph size. This is more realistic than scaling the degree of the nodes as well, as in practical applications, the degree often represents a physical communication limit, which does not increase when more neighbours are in the network.
The obtained results are shown in Table 4. Although the network is only trained on small graphs with $m=8$ nodes, we observe that all learned methods generalize well and outperform the baselines even on the OOD data. The performance of all methods drops compared to the original data shown in Table 1, but does not continue to worsen when increasing the graph size further.
Table 4: Results for OOD generalization with increasing number of nodes in the distributed system after $K=10$ steps. For each size, we solve 2020 problem instances and report the mean performance. All models are trained using problem instances with $m=8$ nodes. Best method is bold, second best underlined.
m=16 m=32 m=64 m=128
Consensus Error Consensus Error Consensus Error Consensus Error Consensus
Fixed 22.92 21.26 18.04 17.24 16.90 16.51 20.07 19.88
Adaptive 25.01 24.00 21.37 20.89 20.56 20.36 24.58 24.48
Global $\alpha$ 18.91 18.16 16.11 15.82 16.30 16.15 19.19 19.13
Local $\alpha_{i}$ 15.90 15.74 13.16 13.04 13.74 13.69 16.37 16.34
Weight $e_{ij}$ 19.63 17.90 17.99 17.29 16.21 15.88 19.76 19.59
Combined 12.93 12.47 10.91 10.58 11.65 11.54 13.95 13.90
Additional ADMM iterations
In order to show the performance of our method when increasing the number of iterations even further beyond the training regime, we showcase the performance for $100$ iterations in Figure 4. Figure 2 showed the same results until $20$ iterations are reached.
The results show that the learned methods continue to outperform the baselines, even far beyond the trained iterations. The only exception to this is the combined method in the least-squares problem, which slows down after 50 iterations. Particularly noteworthy is the strong performance of the method with learned communication matrix, as the learned matrix is used in all iterations and not only in the trained ones. However, using the learned communication matrix together with the learned step sizes performs best in the consensus example while achieving the worst in the least squares problem.
Figure 4: Semi-log plots of the mean relative objective over all $100$ test instances in the two problem classes for $100$ iterations. Note that this is far beyond the trained unrolling of $K=10$ iterations.
Appendix E Algorithms
E.1 Decentralized, distributed ADMM
@@
In the following, we state the decentralized, distributed ADMM algorithm (Algorithm 1) from makh2017admm adapted to our notation. As described in Section 3.2, each iteration updates all three iterates $(x_{i},y_{i},\lambda_{i})$ and includes two communication steps to pass information about the current iterates to the neighbors in $\mathcal{G}$.
@@
E.2 Message-passing formulation of distributed ADMM